OpenStudy (imsleeziboii):

Simplify square root of 18.

OpenStudy (fwizbang):

What numbers multiply together to give 18?

OpenStudy (imsleeziboii):

9 and 2

OpenStudy (imsleeziboii):

6 and 3

OpenStudy (imsleeziboii):

18 and 1

OpenStudy (fwizbang):

good. Is either of these a perfect square?

OpenStudy (imsleeziboii):

what is a perfect square?

OpenStudy (fwizbang):

a number that is the square of another number. So 5^2=25 , making 25 a perfect square.

OpenStudy (imsleeziboii):

neither of those is a perfect square.

OpenStudy (imsleeziboii):

wait

OpenStudy (fwizbang):

you gave me six numbers, 2,9,6,3, 18 and 1. are any of those perfect squares?

OpenStudy (imsleeziboii):

i dont think so?

OpenStudy (fwizbang):

1 x 1 =1 2 x 2 =4 3 x 3 = ?

OpenStudy (imsleeziboii):

9

OpenStudy (fwizbang):

So 9 is a perfect square since its 3 squared.

OpenStudy (imsleeziboii):

Ohh that makes sense

OpenStudy (imsleeziboii):

so 3 to the square root of 2

OpenStudy (fwizbang):

so now if \[\sqrt{18} = \sqrt{9 \times 2} \] what can we say about \[\sqrt{18}?\]

OpenStudy (imsleeziboii):

square root of 18 is 9 to the square root of 2?

OpenStudy (fwizbang):

not quite. Think about the \[\sqrt{36} =\sqrt{4 \times 9} = \sqrt{4} \times \sqrt{9} = 2 x 3 =6\]

OpenStudy (imsleeziboii):

I dont get it

OpenStudy (fwizbang):

If I have two numbers, a and b, its always true that\[\sqrt{a time b} = \sqrt{a} \times \sqrt{b}\], so if 18 = 9 x 2 ....

OpenStudy (imsleeziboii):

\[2\sqrt{?9}\]

OpenStudy (fwizbang):

make that \[\sqrt{a \times b} = \sqrt{a} \times \sqrt{b}\]

OpenStudy (fwizbang):

In oour case, a x b =18, so we can take a=9 and b=2...

OpenStudy (imsleeziboii):

ok

OpenStudy (imsleeziboii):

is that the answer?

OpenStudy (fwizbang):

\[2\sqrt{9}\] isn't quite correct, but you're getting close

OpenStudy (imsleeziboii):

\[9\sqrt{2}\]

OpenStudy (fwizbang):

closer remember that \[\sqrt{a \times b} = \sqrt{a} \times \sqrt{b}\] so you need to take the square root of both factors

OpenStudy (imsleeziboii):

ohh so its \[3\]

OpenStudy (imsleeziboii):

\[3\sqrt{2}\]

OpenStudy (fwizbang):

perfect!

OpenStudy (imsleeziboii):

sweet. ty